Sol Summary – December 15th

MDRS Sol Summary Report for Sol 14

Summary Title: Signing off!

Mission Status:

It’s been a please working with all of you, and thanks for your help.

It was a fitting send off to our team that our day started with a
generator failure. Upon our morning rounds after waking up it was
noticed the Generator had failed with an ‘Underspeed’ Fault. We were
happy to see that SOC was holding steady at 100% (thank you solar
arrays!) but this was a worrying sign, especially as we didn’t want to
hand over a crippled ship to the oncoming crew.

Investigation into the generator showed that the loose connections
that were repaired previously were still holding strong, so there was
no immediate cause of the generator fault. After attempting to
restart the generator a low battery voltage was observed (reading
7.6volts, supposed to be 9.6 volts). It appears that after the
generator failed it drained the battery down displaying the fault
message on its status plate. After a jumpstart from our friendly
neighbors we were able to get the generator back up and running. This
issue is likely to re-occur as we are still not sure of the original
cause. Our next step will be to Remove and Replace the generator
batteries, fingers crossed this fixes the problem. Best of luck to
future crews in continuing to troubleshoot this issue.

Besides that morning excitement the day was one of cleaning and
packing, with a short break in the afternoon to do a short EVA around
the HAB for some last minute filming to get the shots we’ve missed so
far this mission. We hope to leave this place better than we found
it.

This report comes to you with the assistance of Mission Support as our
Internet allowance has run out. Sometime today the same laptop that
killed our bandwidth before had WiFi turned On, after which it sucked
our bandwidth dry again. Still don’t know why this particular laptop
uses bandwidth at such a prodigious rate, but we’ve turned it off
again and comms should be returned to normal.

Hope to work with you all again in the future.

Ad Astra Per Ardua,

Crew 184.

Sol Activity Summary:

1. Filming EVA

2. Cleaning / Packing

4. Reports and general HAB maintenance / checks

6. Generator Failure and Jumpstart

Look Ahead Plan:

1. Crew Handover Training

Anomalies in work:

1. HAB Leak from under EVA door continuing (per Mission Control
leak is expected and has been present for months, presumed to be from
Kitchen piping)

2. Generator Underspeed Fault

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: None today.

Reports to be filed:

Greenhab

Operations

Journalist

Astronomy

Sol Summary

Mission Summary

Support Requested:

1. None

Sol Summary – December 14th

Summary Title:  Wrapping Up!

Mission Status:

Today marks the end of our scientific mission and the completion of our three remaining scientific objectives.

Our martian day sleep study ends tonight with a 4am bedtime, 12pm wake up.  Our team prepped extensively for this and arranged numerous protocols to handle crew fatigue and performance drop off due to tiredness.  Luckily we haven’t had to use any!  Our team has remained in good spirits and been able to keep to our sleep schedule well throughout our mission, with the occasional controlled nap being used to support a flagging crewmember!  We’ve focused on mentally demanding activities and music to keep the crew focused and alert in the evenings to allow us to hit our sleep targets.

Today we successfully completed our last Matryoshka objective!  This was the farthest we’ve had to travel for an EVA, which is why we left it till last to capitalize on the experience we’ve gained in the past two weeks.  We took the Deimos Rover and an ATV to ensure we had sufficient range to meet our objectives and departed from the HAB at around 1pm, arriving at our objective at around 2pm in the Blue Hills area.  Our goal was to get samples from the ridgeline as it washed out into the plane of dirt to try to identify different materials in the mixing area between the cliffs and plains.  We successfully took samples from this area and several isolated hilltops in the area and packed up from or last EVA to return home.

We weren’t going to get off easy however, as that hasn’t been our type of mission.  Due to user error when we started home the parking break was left on for ~ 5 minutes, which resulted in a precipitous power drop off from ~ 80% to 40%.  After the problem was noticed and corrected we discussed options.  The Blue Hills area is actually a fairly close walking distance from the HAB, but a long driving distance, so the decision was weather to leave the rover and walk back to the HAB or to take the long way around on the roads and trust the Rover to make it.  After assessing the fact that the Rover utilized ~ 20% battery to get to this location, had 40% remaining, and that we had a functioning ATV to get help if we broke down on the way we decided we had high confidence in getting the Rover back to the HAB and started the long drive back.  It made for an anxious drive back and frequent map checks to ensure we were going the right way but we managed to successfully coax Deimos back to the HAB just as the low battery light turned on at 20% SOC.  Another successful, and exciting, EVA!

Today we completed our Anesthesia runs ‘for score’.  Under the supervision and evaluation of our resident doctor two crewmembers at a time formed a team and had to evaluate a test mold he has created to mimic a human leg.  Inside the mold the team must use the instructional videos and procedures provided to identify various components inside the mold (veins / arteries / nerves / etc. . .), identify the proper spot to apply anesthesia, then go through the process of delivering anesthesia to the leg.  Hopefully this research and experimentation makes good progress towards developing medical trainers that can be used with astronauts on a long duration mission that may not have anyone on board with medical expertise.

Unfortunately cloud cover prevented any viewing of the Geminid meteor shower last night (of course the only cloudy night was the one with the meteor shower!)  It is currently clear and we intend to try again tonight to see if we can catch the tail end of the meteor shower.

Tomorrow we focus on MDRS cleaning, packing, and transition to the next crew.  We will also be doing an EVA strictly in the environs of MDRS to get some last filming done as requested by our journalist.

Sol Activity Summary:

  1. Matryoshka EVA Site #7a
  2. Exercise
  3. General HAB Cleaning / Maintenance / Reports
  4. Anesthesia Experiment
  5. Beginning of cleanup / packing
  6. (Hopefully)  Telescope / Geminid Observation tonight due to clouds last night!

Look Ahead Plan:

  1. EVA Filming around the HAB
  2. Solar Observation / Observatory Shutdown
  3. MDRS Cleaning
  4. Lessons Learned Compilation
  5. End of Mission Food Inventory
  6. Mission Summary
  7. Exercise
  8. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.    HAB Water Leak continues from under EVA door.  It appears to be getting worse.

2.  Power System Misconfiguration ongoing.

Weather: Warm (above freezing)

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

Sol Summary

Journalist

Support Requested:  None

Sol Summary – December 13th

Summary Title:  In the final stretch (and clouds)

Mission Status:

We are getting close to the end of our mission, and the focus of the crew is maximizing the time we have left to complete our science objectives.

Today in the Matryoshka EVA series. Crewmembers Trisha Randazzo, Joshua Hunt, and Willie Schumann traveled to “Matrysoshka Site 6” which was located in the far north at the begging of Lith Canyon. After a long drive to the site with

We attempted to do solar observation tonight but were unfortunately stymied by extensive cloud cover all day long today.

This evening our team will focus on Anesthesia Experiment runs.  Once complete our team is currently hoping the clouds clear to allow us to watch the Geminid meteor shower which peaks tonight.  Shannon also allowed us to assemble her old telescope for use tonight.  Fingers crossed the clouds clear!

We preparing for our last ‘science day’ on Thursday, as we finish up both Matryoshka, Bricks, and Anesthesia experiments.  Friday is being held in reserve and is also allocated to cleaning the HAB and writing up lessons learned for the next crew.  The end is fast approaching!

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Exercise
  4. General HAB Cleaning / Maintenance / Reports
  5. Telescope Assembly
  6. Anesthesia Experiment Runs this evening
  7. (Hopefully)  Geminid Observation!

Look Ahead Plan:

  1. Matryoshka EVA Site #7a
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.    HAB Leak from under EVA door.  No crew has showered yet today and a significant leak was observed by returning EVA crew.  It is likely this leak is coming from the Kitchen as the dishes were washed immediately prior to the crew returning.

2.  Power System Misconfiguration ongoing.

Weather: Mild and Cloudy

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

Support Requested:

1.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce

Sol Summary – December 12th

Summary Title:  Getting back on routine (and our first EVA failure!).

Mission Status:

Today was a successful day and will hopefully  mark out the routine we will follow until the end of our mission.  The generator has been completely functional now for two days since our repairs over the weekend which is thankful news, our power problems seem to have abated themselves for the mission (though I may be jinxing myself. . . .).  In addition, the weather has been getting warmer which is a welcome relief and allows me to finally stop wearing my thermals, which is a welcome relief as they could sorely use a wash!

We resumed this week’s round of Matryoshka EVAs with a successful visit and sample collection to Site #5.  Just two more sites to go and we will have completed our most important EVA objective of our mission.  We managed to collect several soil and rock samples from what appeared to look like river channels. We saw shiny glinting on the hills too and found some shale like specimens for analysis. All in all a successful EVA! One thing that has come about is the hazard of the strong sunlight in your face reducing visibility, as there is no sun visor on the helmet and sunglasses are not recommended. Luckily it caused no issues on this EVA, but could potentially be dangerous if sunglasses are to remain forbidden and visors are not added to EVA helmets.

During the EVA return our EVA crew encountered a failure with the Deimos Rover.

Deimos left the HAB at SOC 100% and was driven for ~ 45 minutes to the site with no issues.

When returning from EVA site Deimos started experiencing significant mechanical drag that both lowered the achievable rover speed and increased battery drain.  This problem persisted for roughly 30 minutes until the Rover was able to return to base.  From the onset of the problem it was roughly steady until the turn was made off of cowdung road to return to the HAB.  At that time Rover performance precipitously dropped and was able to achieve little more than walking speed.  Steering was also noticed to be stiff during the return from EVA site.

The EVA crew reports that from a driving perspective it felt like the parking break was on, though visual inspection showed this was not actually the case.  The root cause of the issue was not determined though we will do an inspection during tomorrows EVA.  We suspect the issue may be due to debris in Deimos’s drive train.

At the end of the EVA Deimos was at 29% which is unusual given Deimos’s long range and is likely related to the experienced ‘drag’ when driving Deimos.

For the folks left behind today we were primarily occupied with Solar observation which showed numerous impressive features that made observing a treat.  The sun is waking up!  Hopefully it stays active for the next few days and into the next mission.

We are entering the period of our mission where our Martian Sleep Schedule has moved our days ‘out of phase’ with Earth normal.  Tonight we will be going to sleep at 2:40am and waking up at 10:40am.  We have been primarily occupying our nights with cleaning, Bricks, Anesthesia training, and general maintenance to prepare ourselves to fully utilize the next day’s daylight hours.  Our team has adjusted to the schedule well so far with crew taking naps as needed to stay on schedule.

Thanks for your help!

Sol Activity Summary:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Exercise
  5. General HAB Cleaning / Maintenance / Reports
  6. EVA Suit Inspection / Maintenance

Look Ahead Plan:

  1. Matryoshka EVA Site #6
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. Anesthesia Experiment Runs
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports
  7. Deimos Troubleshooting


Anomalies in work:

1.    HAB Leak from under EVA door (not noticeable today)

2.  Power System Misconfiguration ongoing.

3.  Deimos Rover (see above)

Weather: Mild

Crew Physical Status: All crew in good physical health.

EVA: See Above

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

HSO

Support Requested:

1.    ATV Fuel (we are empty)

2.  Burnable / Non Burnable Trash Pickup from Rear Airlock

3.  Permission to replant seedlings into individual planters so herbs have clean soil.  See Greenhab Report for details.

4.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

5.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken
    16. 1 Can Strawberries
    17. 1 Box Granola
    18. 1 Bottle Soy Sauce

Sol Summary – December 11th

Summary Title:  A nominal day!

Mission Status:

Today has been a thankfully nominal day!  Due to our error with comm’s yesterday and not getting our EVA request in on time we stayed at MDRS and concentrated on catching up on HAB maintenance and other activities that have fallen by the wayside in the last few days.

Today we completed our crew training on our ultrasound machine, which enables us to complete our practice runs with the Astronaut trainer developed by Dr. Sczepaniak in the next few days.

We were informed of some likely solar activity by our friendly neighborhood astronomer which led us to perform some highly anticipated solar observation this afternoon.  While we could not detect any prominence we did detect sunspots for the first time.  We are hoping the sun continues in this period of high activity and future solar observation yields some rewarding images.

We continued with our communication experiment and have noticed significant improvement among the crew in terms of numbers of errors and completion times in our building tasks.  We have assembled a lesson’s learned list of communication tips to allow quick and efficient communication among various types of time delay.  We will continue this experiment in various permutations until our mission ends, and then expect to hand over our experiment materials to make this a permanent installation at MDRS that crews can participate in as desired.

We also completed our CPR techniques experiment runs on all crew, analyzing the various ways that more force than body weight can be utilized during a CPR medical emergency.  It was interesting that most crew actually utilized different techniques during this experiment run, with us comparing notes to identify the optimal method.

Our crew has returned to our exercise routine today with all exercise equipment completely assembled.  This has proven to be a crew favorite and we also intend to donate this equipment to MDRS.  We have assembled a stationary exercise bike and a series of medicine balls for weights exercises.

The remainder of the day was filled with HAB maintenance, static water tank transfer, cleaning, some much needed love for our friendly plants in the Greenhab, and assembly of equipment delivered by Mission Control.  We are excited to finish out our mission and eager to get back out into the Martian environment with our 5th Matryoshka experiment site tomorrow.

Thanks for your assistance!

Sol Activity Summary:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports
  7. EVA Suit Inspection / Maintenance
  8. Washvac Assembly
  9. Power system monitoring
  10. Greenhab Planting
  11. Anesthesia Ultrasound Training
  12. Static Water Tank Transfer

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing.  Generator turned off ~ 10am.  SOC stayed close to 100% all day long.

3.  EVA Suit Malfunctions – See Operations Report

Weather: Cold, clear

Crew Physical Status: All crew in good physical health.

EVA: N/A

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Journalist

Astronomy

Support Requested:

1.    ATV Fuel

2.  Burnable / Non Burnable Trash Pickup from Rear Airlock

3.  Permission to replant seedlings into individual planters so herbs have clean soil.  See Greenhab Report for details.

4.  We have an extra CO detector that is marked ‘For Greenhab’.  The current detector in Greenhab is both a Smoke Alarm / CO Sensor.  Please advise whether you would like us to install a redundant CO Sensor in the Greenhab or return it to the HAB and mark it as a replacement part.

3.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken

Sol Summary – December 10th

MDRS Daily Summary Report for Sol 9

Summary Title:  Return of the Sim!

Mission Status:

Yesterday and today has been been primarily ‘out of sim’ due to the numerous HAB failures, including our evacuation last night due to a loss of power.  We spent the morning today working to repair the generator and appear to have gotten it in working order, though further work is necessary to ensure it’s long term successful operation (see Shannon’s email for details).  The solar array system is still misconfigured so power problems may yet pop up during the rest of our sim but we will take them as we go and troubleshoot, as a real martian crew would.

This afternoon we took the rovers out to drain the batteries per Mission Control request, once returned to the HAB we re-entered SIM at roughly 5pm.

It’s been a rough first week, but we are looking forward to getting back into sim and finishing out our mission successfully.

Sol Activity Summary:

1.      Out of SIM maintenance work

2.     General HAB cleaning / reports

Look Ahead Plan:

  1. Matryoshka EVA Site #5
  2. ‘Bricks’ communication experiment
  3. Solar Observation
  4. CPR Techniques
  5. Exercise
  6. General HAB Cleaning / Maintenance / Reports


Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration ongoing

3.  2 EVA suit fans were non-functional during HAB check today, which may be related to power issues and resulting in them not being charged.  We have connected the suits to power and will assess tomorrow.  

Weather: Cold, but temps are above freezing in mid-day

Crew Physical Status: All crew in good physical health.

EVA: N/A

Reports to be filed:

Greenhab

Operations

EVA Request

Sol Summary

Support Requested:

1.      ATV Fuel

2.     Food Re-Supply

    1. 1 Box Raisin Bran
    2. 1 Box Powdered Milk
    3. 24 packs Instant Ramen
    4. 3 Boxes Pasta (any type)
    5. 2 Box Granola Bars
    6. 1 Box Saltine Cracker
    7. 1 Box Cracker (any type)
    8. 1 Box Hot Chocolate
    9. 1 Box Red Lobster Biscuits
    10. 1 Box Earl Grey Tea
    11. 1 Pack Chocolate Chips
    12. 1 Pack Tortillas
    13. 1 Pack Sour Cream\
    14. 1 Can SPAM
    15. 2 Can Chicken

Sol Summary – December 7th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Back on schedule!

Mission Status:  The past two days have been hectic due to getting in to the swing of things on Mars and the numerous failures we have encountered.  As a consequence, our schedule fell by the wayside and we were largely winging it.  Yesterday we sat down, came to a stable configuration, discussed our lessons learned from our stay so far, and prepared to jump back onto our schedule.

We are still dealing with our power system failure, but today has been a day of focus where our team is efficiently moving from task to task and completing our assigned activities.

We’ve learned today that the low battery SOC we’ve been seeing since our mission started is likely due to a misconfigured battery panel.  Mission Control is seeking help to reprogram it and return our power system to nominal.  In the meantime we will continue to minimize power usage.  Due to SOC ~ 35% it is possible we lose power tonight, our team has prepared for this eventuality and everyone is prepared to handle the cold if necessary, and all team members have light sources on hand overnight.  We have also filled reserve containers with water due the expectation that if power is lost pipes will freeze overnight.

  1. Power system is still smart enough to accept power from generator once SOC starts dropping overnight.  System Nominal!
  2. Power system swaps to generator power.  Generator Failure overnight due to the cold.  Expected loss of power overnight
  3. Power System misconfiguration prevents accepting generator power.  Expected loss of power overnight.

In the meantime, activities!

This morning we have completed Anesthesia training for all crewmembers.  On a long duration mission it is critical crews know how to treat any injuries or other problems that arise, as part of this devising methods of training for crew on a long duration mission is critical.  Our Dr. has devised a mold that simulates arteries / nerves / and veins in a leg that allows crew to practice using ultrasound equipment to identify and isolate specific body parts.  With the training complete crew will be able to complete trial runs where we see how effective this is a training device and how crew performance increases over time.

We also completed our third Matryoshka EVA to SIte #3.  We continue to collect geological samples from our target sites, though this is the first day that our crew is starting to feel the fatigue from EVA, which is hard work!.  Today’s site was exceptionally rocky which gave a very ‘martian’ feel.

Our crew is looking forward to the weekend which we have schedule as ‘light duty’ days to allow ourselves to recuperate from a very busy week!

Sol Activity Summary:

1.      Matryoshka EVA Site #3

2.      Observatory Operation

3.      Anesthesia Training

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel and submission to

6.      Exercise

7.  Brick Experiment

8.  Power System Monitoring

9.  Receipt and stowage of supply drop

10.  CPR techniques (ME-3)

11.  Media Visit Cancelled

Look Ahead Plan:

  1. Matryoshka EVA Site #4
  2. ANAM Winscat Baseline
  3. Musk Observation
  4. CPR Tecniques
  5. Brick Experiment
  6. Filming and Interviews
  7. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door continuing (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.  Power System Misconfiguration

Weather: Still cold (below freezing).

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock (Non Burnable)

2.      Replacement ATV Fuel

Sol Summary – December 6th

Crew 184 Daily Summary Report  06 Dec 2017

MDRS Daily Summary Report for Sol 5

Summary Title:  Continuing failures, schedule replans, and a successful EVA!

Mission Status:  We are still waiting for a nominal morning!  To fix yesterday’s bandwidth problems we powered off all devices in the HAB, we then connected our ‘internet’ laptop to the modem and watched 100MB of bandwidth disappear in ~ 5 minutes.  We concluded there was a background process of some sort operating on that computer as no known programs were running.  We were able to fix this issue by swapping out this laptop and powering down the fault one.  Communication restored!

We exulted in this victory for approximately 10 minutes, then noticed that our generator had failed again and we were back to a powerdown configuration.  Our SOC was holding at 47%, but it was again clear that we were operating with a degraded power system.  Communication with Mission Control has tentatively identified the problem as due to a ‘summer fuel’ blend which has been noted before, which is separate from the oil leak we observed earlier.  The generator will likely continue to be faulty for the next few days until we can procure a ‘winter’ blend.  Winters on Mars can be tough!.  We are continuing to conserve power wherever possible and using the generator minimally to supplement the solar arrays when required.

We resumed nominal operations this afternoon with two primary activities

We decided to finish up the EVA that we left incomplete yesterday, going to Matryoshka Site #2 which is adjacent to the MDRS HAB.  Our EVA prep procedure continues to be challenging, but we are getting better.  We have allocated the best radios and headsets for EVA (recently delivered!) and placed them adjacent to the EVA prep room.  We have also pre-staged all EVA equipment (boots / gloves / first aid kid / emergency water supply / maps / etc . . . ) in the EVA room to minimize the chance that our prebreathe is interrupted by having to go out to the HAB to get supplies.  We have also started carrying backup radios with us since we have had to cut an EVA short due to radios running out of batteries.  The EVA training continues!

Once outside we successfully found samples from numerous geological strata in our target era and bagged them for later delivery and chemical analysis.  We had some excess time upon returning so we also took some of the PR images requested and will uplink them as part of the journalist report.

When the majority of the crew got back to the habitat we did an observatory training sessions where the crew astronomy gave a tutorial in observatory and telescope operation.  There’s something about being on Mars that gives observatory operations extra gravitas as you are already in the ‘space’ mindset.

Sol Activity Summary:

1.      Matryoshka EVA Site #2

2.      Observatory Training Session (no observing)

3.      HAB Cleaning

4.      Reports and general HAB maintenance / checks

5.      Food inventory transfer to excel

6.      Exercise

Look Ahead Plan:

  1. Matryoshka EVA Site #3
  2. Nerve Block Training
  3. CBS News Visit (unconfirmed – request confirmation on whether media crew is coming from Mission Support)
  4. Filming and Interviews
  5. Solar Observation
  6. Communication Experiment (Time Delay)
  7. CPR Technique Training
  8. Exercise / General HAB Maintenance

Anomalies in work:

1.      HAB Leak from under EVA door (per Mission Control leak is expected and has been present for months, presumed to be from Kitchen piping)

2.      Generator failure

3.      Internet bandwidth problem was resolved today.

Weather: Still cold (below freezing), slightly windy

Crew Physical Status: All crew in good physical health.

EVA: See above for EVA details

Reports to be filed:

Greenhab

Operations

EVA Request

Journalist

Astronomy

Sol Summary

Support Requested:

1.      Trash pickup from rear airlock

2.      Replacement ATV Fuel

3.      Additional dish soap

4.  Confirmation of whether media visits will occur

Sol Summary – December 5th

Crew 184 Sol Summary Report  05 Dec 2017

MDRS Sol Summary Report for Sol 4

Summary Title:  Our first failure and major replan

Mission Status:  Today was a hectic day with three major failures our
team dealt with.
1.    Internet – Between 8:30am and 9:30am 800MB of bandwidth was
utilized. It is unknown what caused this but likely culprit was a
device being left on and doing a software update in the background.
Internet communication was lost at this time.  These reports are being
delivered by Shannon today.
2.    Water – During our morning fill of the loft tank the pump
was non-functional.  Possible causes were quickly narrowed down to
either a frozen pipe due to the below freezing temperatures or a
clogged filter. Our team immediately switched to using bottled water
which we keep as an emergency reserve while we troubleshot the issue.
At 1pm we called Mission Control for advice who recommended we connect
the pipe heating system and wait till midday, this resolved the issue
around 3pm and water was restored to the HAB
3.    Power – During our morning check of SOC it was noted that
the generator was not running, having failed sometime during the
night, and that SOC was at 3%.  We immediately instituted a shutdown
of MDRS powered equipment to preserve what little power we had and
investigated the generator.  It was noted that a significant oil leak
had developed from the oil cap, upon tightening the cap the leak
stopped and the generator was restarted.  When returning to the
science dome it was observed that SOC was hovering around 10% and did
not appear to be charging.  After consultation with Mission Control it
was discovered that the generator was in idle due to the power system
not requesting it to go to full delivered power given the fact that
the solar array system was charging the battiers (however slowly) .
In addition the SOC was likely incorrect due to it being a calculated
value and us confusing it due to going off scale.  We will assess SOC
tonight after a full day in the sun and decide forward plan.

Our morning was mostly consumed with maintenance and troubleshooting,
we quickly discovered that when receiving failures our level of
knowledge of the systems was exhausted in how to troubleshoot,
recommend future crew get access to and study manuals for MDRS
equipment prior to stay.

In the afternoon when we got our facilities in a stable configuratrion
four members of our team went on our first Matryoshka EVA. We were
able to go to the Kissing Camel ridge and collect two samples, however
we were unable to go to the other site in URC South due to delays in
getting out since we had to fix our non-nominal systems.

The two sample we collected were to highlight interesting
geomorphological features to be used in outreach activities back in
Oxford. The strike and dip of the ridge component the first sample was
extracted from was also taken. The intention of sample collection and
strike and dip measurement is to determine whether satellite and
outcrop scale imagery data is reliable at predicting these.

Sol Activity Summary:
1.    Troubleshooting and EVA as discussed above
2.    Exercise
3.    Assembling of exercise equipment using martian dirt
(medicine ball complete!)
4.    HAB Cleaning
5.    Solar observation

Look Ahead Plan:
With internet down we are unable to contact our scheduling team.  We
will receive the schedule in the morning when internet bandwidth
restarts and go over it as a team to plan our day.

Anomalies in work:
1.    Yamaha 250 ATV #3 wheel deemed non-recoverable.  We have
parked it off to the side for mission support.  Photos of damage
attached.

Weather: Very cold!  It was below freezing for most of the day.

Crew Physical Status: All in good health.

EVA: Delayed due to non-nominal systems, so we are requesting the same
EVA from Sol 3 but just for site A (mound on URC South) to finish off
the sample collection. The plan and crew utilised will be the exact
same as today.

Reports to be file:
Greenhab
Operations
Astronomy
EVA Request
Journalist
Sol Summary

Support Requested:
Trash placed in the Rear Airlock for pickup.  Burnable trash on the
side with the box.  Non-burnable trash on the other side.

Sol Summary – December 4th

MDRS Daily Summary Report for Sol 3

Summary Title: First full up day of activities!

Mission Status: Today really felt like our first day on mars. It was our first full activity day as well as our first day on EVA! Surprisingly it was consensus among our crew that our first full day was actually the easiest day we’ve had so far! I think the excitement of going outside and our team getting into the swing of things has really made us more efficient.

The primary activity for the day was an EVA on all crew to familiarize ourselves with EVA procedures, suit functionality, and operation of the rovers. Many lessons were learned, and it was good that we got the kinks out on a EVA solely suited to it. This will allow our scientific EVAs to be successful later on, as well as ensure that we can safely perform EVAs at a farther distance from the habitat.

Our team is following a specific protocol for depress / repress operations modeled after ISS EVAs. Our team is doing a 25 minute prebreathe in the EVA Prep room to prevent decompression sickness, once complete we are doing a 5 minute depressurization in the actual airlock. During repress we are allocating 5 minutes for repressurization options, though we have the capability to do an ’emergency repress’ if needed at the risk of rupturing crew eardrums.

Below are the several lessons learned we had during EVA.

  1. Peripheral vision is nearly non-existent when wearing the EVA helmets. This makes it difficult to travel together in an ATV convoy as it is impossible to turn around and check on the status of those behind you. We attempted to mitigate this by placing the radios in VOX and keeping in constant voice contact to maintain proper spacing, we also played ‘leapfrog’ with the ATV’s where the leader would stop after several minutes of driving to allow the two behind him to catch up and pass, ensuring no-one was left behind.
  2. We had several communication failures during EVA due to earpieces falling out of ears. We addressed this via using medical tape to secure the earpieces in place.
  3. We had to abort a prebreathe operation due to forgetting to bring supplies into the EVA Prep room. We instituted a supply check prior to closing the Prep Room door. If necessary to break EVA prebreathe protocol we do a 3 to 1 buy back, i.e. if you break prebreathe for 5 minutes, you add 15 minutes to prebreathe to compensate.
  4. Several of our crew members are trained scuba divers. We found that it was difficult to hear other EVA crew while driving the ATV due to engine and wind noise, we instituted a series of hand signals to allow non-verbal communication which quickly and concisely allowed EVA crew status to be conveyed.
  5. Do not wear high pony tails – recommend low braid. A high braid interferes with the way that your head can move within the EVA helmet and vision is confined to roughly a 45 degree angle down and lower.

During the high winds today we heard a bang as the outer airlock door was blown open. The inner airlock door held closed and thus we didn’t all die! We immediately closed the door to the EVA Prep room to form a second seal to the martian atmosphere. During tomorrows familiarization EVA we will enter the EVA Prep room, close the EVA Prep room door, and use that as a depressurization chamber. Immediately upon departure on tomorrows EVA we will repair the outer airlock door.

We also completed our first scientific objectives today, covered below.

Reduced Gravity CPR Experiment: When on Mars there is a possible complication when performing CPR. With your body weight reduced the typical CPR techniques are non functional due to reduced force on the recipient. Our team is practicing techniques under the supervision of our medical officer to allow force application in excess of body weight and measuring output using a springed scaled.

Communication Delay ‘Bricks’ Experiment: Our first two experiment runs of the brick experiment were performed. We started with a 0 time delay baseline, and increased to a 9 minute time delay with verbal information transfer (martian time delay). Our runs will continue to collect more data and determine the most efficient methods of information transfer.

We also completed some major maintenance tasks today

The Musk Observatory has been set up and put through it’s paces and is ready for operation.

Trailer Tank to Static Water Tank transfer operation – Crew completed the transfer this morning completely filling our static water tank. Heartfelt thanks to the Team supporting us for dropping the trailer tank and instructions on how to transfer the water. The MDRS Hab Manual was a roaring success in allowing us to complete this task!

Sol Activity Summary:

  1. Tasks mentioned above
  2. Exercise
  3. Dirt dehydration
  4. Science Dome organization and clean up
  5. Repair of Roof Hatch hole
  6. Our first shower! Heartily enjoying the new heater
  7. Completion of reports (soon)
  1. EVA Request
  2. Sol Summary
  3. Greenhab
  4. Journalist
  5. Operations
  6. Astronomer
  7. Inventory

WiFi range survey Radio check with all radios (all functional except for 1) Look Ahead Plan:

1. Matryoshka EVA #1

2. Communication ‘Lego Bricks’ experiment

3. Sleep Study / crew medical questionnaires

4. Musk Observatory Observation

5. Journalist interviews and filming

6. Medical status surveys with Crew Doctor throughout the day.

7. CPR Experiment

8. HAB Cleaning

9. Radio Range EVA – This activity is cancelled due to 1 EVA per day limit. It will be rescheduled for a future day. We will work with schedulers to ensure they understand this limit for the future.

Anomalies in work:

  1. ATV wobbly wheel
  2. Non Functional Power Strip (removed and replaced)
  3. 3 Non Functional Radios (does not impact operations)
  4. HAB paneling peeling off due to wind damag

Weather: Very mild weather. This was welcome after yesterday.

Crew Physical Status: All crew in good health.

EVA: 2 Familarization EVAs today

Support Requested:

  1. ATV Fuel
  2. Food Drop (not urgent)

.